OpenStudy (anonymous):

please help me solve! 9(8 - x) = 3x.

OpenStudy (anonymous):

So can you expand 9(8-x) = ?

OpenStudy (anonymous):

72-9x?

OpenStudy (anonymous):

Yes. 72 - 9x = 3x. So now if you shift 9x to the other side, what would the equation become?

OpenStudy (anonymous):

you would divide and that would equal 3x?

OpenStudy (anonymous):



OpenStudy (anonymous):

actually nope. 72 - 9x = 3x. just shift -9x to other side it will become +9x 72 = 3x+9x 72 = 12x. solve for x.

OpenStudy (anonymous):

x=6

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

i get it